OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 1) and (–1, –5)?

OpenStudy (tester97):

Do you know the formula for finding slope?

OpenStudy (anonymous):

NOOOOO

OpenStudy (anonymous):

(X1 X2) Y1,Y2

OpenStudy (tester97):

Well the formula is this: \[\frac{ y^2-y^1 }{ x^2-x^1 }\] Can you fill in the variables youself??

OpenStudy (tester97):

*yourself Note: If not then i can help you

OpenStudy (anonymous):

1,-1/1,-5

OpenStudy (tester97):

No you have to have the x variables on the bottom and the y on the top like this. \[\frac{ 1-(-5) }{ 1-(-1) }\] But do you notice you have to subtract a negative?

OpenStudy (anonymous):

YEAA

OpenStudy (anonymous):

6,2

OpenStudy (tester97):

Well when that happens that means that you basically should just add the two. So yes its going to be \[\frac{ 6 }{ 2 }=3\] So your slope is 3 ^_^